QUESTION 1

  1. The information in this question pertains to this question and the next four questions.

    Bama Corp. has sold British pound call options for specula­tive purposes. The option premium was $.06 per unit, and the exercise price was $1.58. Bama will purchase the pounds on the day the options are exercised (if the options are exer­cised) in order to fulfill its obligation. For each possible spot rate, find in the net profit (or loss) to Bama Corp. per unit of currency, assuming the listed spot rate exists at the time the purchaser of the call options considers exercising them.

    Spot rate = $1.60. Net profit or loss is:

1.00000 points   

QUESTION 2

  1. Spot rate is $1.62. Net profit or loss per unit is:

1.00000 points   

QUESTION 3

  1. Spot rate is $1.64. Net profit or loss per unit is:
Answer #1

Call Option Premium = $ 0.06

Exercise Price= $1.58

Question1) Spot Rate = $1.60

Spot rate greater than Exercise price so the call option will be exercised.

Net profit by Bama corp = Premium Received - loss from execution of call option = $0.06 - ($1.6-$1.58) = $0.04

Question2) Spot Rate = $1.62

Spot rate greater than Exercise price so the call option will be exercised.

Net profit by Bama corp = Premium Received - loss from execution of call option = $0.06 - ($1.62-$1.58) = $0.02

Question3) Spot Rate = $1.64

Spot rate greater than Exercise price so the call option will be exercised.

Net profit/Loss by Bama corp = Premium Received - loss from execution of call option = $0.06 - ($1.64-$1.58) = $0